OpenStudy (lcnettech):

-6 times (-4.2) divided by (-3) =?

OpenStudy (lcnettech):

please help urgent i have to go to bed soon . and it's already 11:29pm

OpenStudy (anonymous):

-6 times -4.2 o-o

OpenStudy (anonymous):

= 10.2

OpenStudy (lcnettech):

negative 6 times negative 4.2 divided by negative 3

OpenStudy (anonymous):

divided by negative 3 = -13.2

OpenStudy (anonymous):

/o.o\

OpenStudy (anonymous):

i think

OpenStudy (lcnettech):

ok thanks

OpenStudy (anonymous):

ty :P

OpenStudy (lcnettech):

wait these are the snswer choices

OpenStudy (anonymous):

its 8.4

OpenStudy (kropot72):

First calculate the value of the numerator. Then divide the result by negative 3.\[\large \frac{-6\times-4.2}{-3}=?\]

OpenStudy (lcnettech):

is he correct kropot72?

OpenStudy (kropot72):

No.

OpenStudy (lcnettech):

so it's 2.1 because negative divided by a negative is always positive.

OpenStudy (lcnettech):

am i right?

OpenStudy (anonymous):

its -8.4

OpenStudy (kropot72):

The result of multiplying out the numerator is a positive number. A positive divided by a negative is a negative.

OpenStudy (lcnettech):

no because negative divided by negative is ALWAYS positive.

OpenStudy (lcnettech):

@keeks4481

OpenStudy (lcnettech):

so its between c and d

OpenStudy (lcnettech):

is it C?

OpenStudy (anonymous):

-6*(-4.2)= 25.2 divide by -3 is -8.4

OpenStudy (anonymous):

its A

OpenStudy (lcnettech):

o right i forgot since -6 * -4.2 = 25.2 then divided by 3 = -8.4 thanks

OpenStudy (anonymous):

No problem

OpenStudy (lcnettech):

sorry i just got a little confussed i was thinking about division not multiplication . again sorry.

OpenStudy (lcnettech):

ok bye thanks for your help every one...............................

OpenStudy (anonymous):

its cool